# vHeat: Building Vision Models upon Heat Conduction

### v1

@@ 4.3 Visualization Analysis
Visual Heat Conduction. In Fig. 5, we visualize the temperature $U^{t}$ defined in Eq. (10) under predicted $k$ when a random patch is taken as the heat source. With a predicted $k$, vHeat delivers self-adaptive visual heat conduction. With the increase of heat conduction time ($t$), the correlation of the selected patch to the whole image is enhanced. Please refer to Sec. F in Appendix for more visualization instances.
Figure 6: Visualization of the effective receptive fields (ERF) [36]. Pixels of higher intensity indicate larger responses with the central pixel.
Receptive Field. The Effective Receptive Field (ERF) [36] of an output unit denotes the region of input that contains elements with a non-negligible influence on that unit. In Fig. 6, ResNet, ConNeXT, and Swin have local ERF. DeiT [52], HiViT [65], and vHeat exhibit global ERFs. The difference lies in that DeiT and HiViT have a $\mathcal{O}(N^{2})$ complexity while vHeat enjoys $\mathcal{O}(N^{1.5})$ complexity.
Figure 7: Left / Mid / Right: Throughput / GPU memory / FLOPs under different image resolutions. The throughput and GPU memory are tested on 80 GB Tesla A100 GPUs with batch size 64. Swin-B is tested with scaled window size here.
4.4 Computational Cost
The comparisons of throughput / GPU memory / FLOPs of vHeat-B and other ViTs are shown in Fig. 7. Thanks to HCO’s $\mathcal{O}(N^{1.5})$ computational complexity $w.r.t.$ $N$ image patches, vHeat-B has a significant superiority over other base-level ViT models on throughput / FLOPs. Fig. 7 (right) shows that with the increase of input image resolution, vHeat enjoys the slowest increase of computational overhead. Fig. 7 (middle) shows that vHeat requires 70% GPU memory less than Swin-Transformer given large input images. Given the larger image resolution, the superiority becomes larger. These demonstrate vHeat’s great potential to handle high-resolution images.
4.5 Ablation of Thermal Diffusivity
Table 3: Evaluating thermal diffusivity $k$.
Settings Acc
@@ 4.5 Ablation of Thermal Diffusivity
To show the effectiveness of shared FVEs, we conduct the following experiments on ImageNet-1K. (1) Fix the thermal diffusivity $k=0.0/1.0/10.0$. (2) Treat $k$ as a learnable parameter for each layer. (3) Use individual FVEs to predict $k$ for each layer. As shown in Table  3, when $k=0.0$, the visual heat conduction doesn’t work. A larger fixed $k$ value, $e.g.$, $k=5.0$, enables HCO to work isotropically without considering the image content and the performance reaches $81.7\%$ top-1 accuracy. Predicting $k$ by FVEs outperforms treating $k$ as a learnable parameter, which may be attributed to the strengthened prior knowledge of frequency values provided by FVEs. Please refer to Sec. E.5 in Appendix for the detailed analysis. When $k$ is predicted by shared FVEs, the performance improves to 82.2%, which validates shared FVEs can effectively reduce the learning diffusivity and further improve the performance.
@@ 5 Conclusion
We propose vHeat, a visual backbone model that leverages the advantages of global receptive fields, low complexity, and interoperability. The effectiveness of the vHeat model family, including vHeat-T/S/B models, has been demonstrated through extensive experiments and ablation studies, significantly outperforming popular CNNs and ViTs. The results highlight the potential of vHeat as a new paradigm for vision representation learning, offering fresh insights for the development of physics-inspired vision models.
Limitations. Based on the physical heat conduction, the learning process of vHeat may become challenging when distant information conduction is required, as it needs extensive training to effectively perceive long-range dependencies. Moreover, masked image modeling serves as an effective self-supervised learning paradigm for ViTs. By now, we have not yet developed a self-supervised learning method for vHeat, which is left for the future work.
@@ Appendix A Motivation
Appendix B HCO implementation using $\mathbf{DCT_{2D}}$ and $\mathbf{IDCT_{2D}}$
@@ Appendix C Experimental Settings
Model configurations. The configurations of vHeat-T/S/B models are shown in Table  4.
@@
Layers (2, 2, 6, 2) (2, 2, 18, 2) (2, 2, 18, 2)
Channels (96, 192, 384, 768) (96, 192, 384, 768) (128, 256, 512, 1024)
Image Classification. Following the standard evaluation protocol used in [31], all vHeat series are trained from scratch for $300$ epochs and warmed up for the first $20$ epochs. We utilize the AdamW optimizer [34] during the training process with betas set to $(0.9,0.999)$, a momentum of $0.9$, a cosine decay learning rate scheduler, an initial learning rate of $2\times 10^{-3}$, a weight decay of $0.08$, and a batch size of $2048$. The drop path rates are set to $0.1/0.3/0.5$ for vHeat-T/S/B, respectively. Other techniques such as label smoothing ($0.1$) and exponential moving average (EMA) are also applied. No further training techniques are employed beyond these for a fair comparison. The training of vHeat-T/S/B takes 5/7/8.5 minutes per epoch on Tesla 16$\times$×V100 GPUs.
@@
Appendix D Performance Comparison
The complete comparison of vHeat and other vision models on ImageNet-1K is shown in Table  5.
Table 5: Performance comparison of image classification on ImageNet-1K.
Method
Image
size
#Param. FLOPs
Test Throughput
(img/s)
ImageNet
top-1 acc. (%)
DeiT-S [52] 2242 22M 4.6G 1761 79.8
DeiT-B [52] 2242 86M 17.5G 503 81.8
ConvNeXt-T [33] 2242 29M 4.5G 1198 82.1
ConvNeXt-S [33] 2242 50M 8.7G 684 83.1
ConvNeXt-B [33] 2242 89M 15.4G 436 83.8
HiViT-T [65] 2242 19M 4.6G 1393 82.1
HiViT-S [65] 2242 38M 9.1G 712 83.5
HiViT-B [65] 2242 66M 15.9G 456 83.8
XCiT-S/12 [1] 2242 26M 4.8G 1283 82.0
XCiT-S/24 [1] 2242 48M 9.1G 671 82.6
XCiT-M/24 [1] 2242 84M 16.2G 423 82.7
Vim-S [67] 2242 26M - 811 80.5
Swin-T [32] 2242 28M 4.6G 1244 81.3
Swin-S [32] 2242 50M 8.7G 718 83.0
Swin-B [32] 2242 88M 15.4G 458 83.5
vHeat-T 2242 29M 4.6G 1514 82.2
vHeat-S 2242 50M 8.5G 945 83.6
vHeat-B 2242 87M 14.9G 661 83.9
Appendix E Additional Ablation Studies
E.1 Interpolation of FVEs/$k$ for downstream tasks
We have tried several approaches to align the shape for ablation. (1) Directly interpolate FVEs to the target shape of the input image. (2) Add 0 to the lower right region of FVEs to align the target shape. (3) Add 0 to the lower right region of FVEs to $512\times 512$, and interpolate to the target shape. (4) Directly interpolate the predicted thermal diffusivity $k$ to the target shape. The results are summarized in Table  6. Through the comparison, we select adding 0, then interpolating FVEs to the target shape for all downstream tasks.
@@ E.1 Interpolation of FVEs/$k$ for downstream tasks
E.2 Plain vHeat model
We’ve tested the performance of plain vHeat-B on ImageNet-1K classification. Keeping the same as DeiT-B, plain vHeat-B has 12 HCO layers, 768 embedding channels and the patch size is set to 16. Results are shown in Table  7. The superiority of plain vHeat-B over DeiT-B also validates the effectiveness of vHeat model.
Table 7: Evaluating different methods to align the shape of FVEs when loading ImageNet-1K pre-trained vHeat-B weights for detection and segmentation on COCO.
@@ E.2 Plain vHeat model
E.3 Depth-wise convolution
We conduct experiments to validate the performance improvement from DWConv. We replace depth-wise convolution with layer normalization for vHeat-B. Results are summarized in Table  8, and vHeat-B achieves 83.7% Top-1 accuracy on ImageNet-1K classification, 0.2% lower than with DWConv, which validates the main gains come from the proposed HCO. Besides, when $k$ is fixed as a large value, e.g. $k=10.0$, replacing DWConv with layer normalization causes a significant performance drop (-0.7% top-1 accuracy). The comparison validates predicting $k$ by FVEs can effectively improve the robustness of vHeat.
Additionally, we train vHeat without DWConv with a different recipe from vHeat with DWConv. The batch size is set as 1024, the initial learning rate is set as $1\times 10^{-3}$, and the weight decay is set as 0.05.
@@ E.3 Depth-wise convolution
vHeat-B 83.9
vHeat-B 83.7 (-0.2)
vHeat-B (fix $k$=10.0) 83.5
vHeat-B (fix $k$=10.0) 82.8 (-0.7)
E.4 Global filters
Considering HCO works as a global filter in the frequency domain for visual heat conduction, we compare vHeat with (1) GFNet [41], and (2) replacing HCO with the operators proposed in GFNet for ablation. Results are summarized in Table  9, vHeat-S has a large superiority over GFNet-H-B under approximate model scale. Besides, replacing HCO with operations proposed in GFNet achieves lower performance, which validates the effectiveness of the proposed HCO and visual heat conduction modeling for representation.
Table 9: Comparison of vHeat with global filters, where vHeat-B denotes replacing HCO with operators proposed in GFNet.
Model #Param. FLOPs Acc
vHeat-S 50M 8.5G 83.6
GFNet-H-B [41] 54M 8.4G 82.9
vHeat-B 87M 14.9G 83.9
vHeat-B 87M 14.9G 83.5
E.5 Predicting $k$ by FVEs vs. treating $k$ as a learnable parameter
After performing DCT, the features lack explicit frequency value, while FVEs provide the model with prior knowledge of frequency values. Similar to how the introduction of positional encoding can enhance performance even in models that include positional information [17], predicting $k$ by FVEs, rather than treating $k$ as a learnable parameter, reinforces prior frequency information and more clearly represents the relationship between frequency and thermal diffusivity.
@@ Appendix F Heat Conduction Visualization
Appendix G Feature Map Visualization
We visualize the feature before/after HCO in a random layer in stage 2 with randomly selected images as input, Fig.  9. Before HCO, only a few regions of the foreground object are activated. After HCO, almost the entire foreground object is activated intensively.
@@ Appendix G Feature Map Visualization
Appendix H Analysis of $k$ in each layer
We calculate average values of $k$ in each layer of ImageNet-1K classification pre-trained vHeat-Tiny, Fig.  10. In stage 2 and stage 3, average values of $k$ corresponding to deeper layers are larger, indicating that the visual heat conduction effect of deeper layers is stronger, leading to faster and farther overall content propagation.
Figure 10: Average values of $k$ in each layer.

### v2

@@ 4.2 Analysis of Dynamic Locality
Visual Heat Conduction. The proposed vHeat works upon an adaptive filtering mechanism. To verify this claim, in Fig. 6, we visualize the temperature $U^{t}$ defined in  (9) under predicted $k$ when a random patch is taken as the heat source. With a predicted $k$, vHeat delivers self-adaptive visual heat conduction. As the heat conduction time ($t$) increases, the correlation between the selected patch and the entire image improves, which effectively filters out unrelated patches in the frequency domain. Please refer to Sec. E in the supplementary for vHeat’s effective receptive field visualization.
Figure 6: Temperature distribution ($U^{t}$) when using a randomly selected patch as the heat source. Please zoom in for details.
Ablation of thermal diffusivity. To demonstrate the effectiveness of shared FVEs, we conduct the following experiments on ImageNet-1K: (1) Fix the thermal diffusivity $k=0.0/1.0/10.0$, (2) Treat $k$ as a learnable parameter for each layer, and (3) Use individual FVEs to predict $k$ for each layer. As shown in Table 6, when $k=0.0$, the visual heat conduction does not function effectively. A larger fixed $k$ value, e.g., $k=5.0$, allows HCO to operate isotropically without considering the image content, achieving a top-1 accuracy of 81.7%. Predicting $k$ via FVEs outperforms treating $k$ as a learnable parameter, likely due to the enhanced prior knowledge of frequency values provided by FVEs. After performing DCT, the features lose explicit frequency values, whereas FVEs offer the model prior frequency information. Similar to how positional encoding enhances performance in models that already incorporate positional information [19], predicting $k$ using FVEs rather than treating it as a learnable parameter reinforces the frequency prior and clarifies the relationship between frequency and thermal diffusivity. When $k$ is predicted by shared FVEs, the performance improves to 82.2%, validating that shared FVEs effectively reduce learning diffusivity and further enhance performance.
Table 6: Evaluating thermal diffusivity $k$ with vHeat-T.
Settings top-1 acc. (%)
@@
Table 7: Comparison of vHeat with global filters, where vHeat-B denotes replacing HCOs in vHeat-B with GFNet operators.
Model #Param. FLOPs top-1 acc. (%)
GFNet-H-B 54M 8.4G 82.9
vHeat-S 50M 8.5G 83.6
vHeat-B 68M 11.2G 83.5
vHeat-B 68M 11.2G 84.0
4.3 Comparison With Global Filters
To simulate physical heat conduction in vision tasks, we developed the Heat Conduction Operator (HCO) in the frequency domain. We compare HCO with (1) GFNet [47], a model using global filters in the frequency domain, and (2) an ablation where HCOs are replaced by GFNet’s frequency-domain operators. The results in Table 7 show that vHeat-S outperforms GFNet-H-B with a similar model size. Furthermore, when HCOs are substituted with GFNet’s operators, there is a noticeable performance drop, underscoring the unique advantages of the proposed HCO. These findings validate the efficacy of our proposed HCO and the underlying visual heat conduction modeling in enhancing feature representation.
@@ 5 Conclusion
We introduce vHeat, a visual representation model that integrates the advantages of global receptive fields, computational efficiency, and enhanced interpretability. Extensive experiments and ablation studies demonstrate the efficiency and effectiveness of the vHeat model family, including vHeat-T/S/B models, which significantly outperform popular CNNs and ViTs. These results highlight vHeat’s potential as a novel paradigm for vision representation learning, offering fresh insights for the development of physics-inspired representation models in computer vision.
6 Acknowledgement
This work was supported by National Natural Science Foundation of China (NSFC) under Grant 62225208 and 62450046 and CAS Project for CAS Project for Young Scientists in Basic Research under Grant No.YSBR-117.
@@ Appendix A Motivation
Appendix B HCO implementation using $\text{DCT}_{\text{2D}}$ and $\text{IDCT}_{\text{2D}}$
@@
We compared our implementation of DCT/IDCT in vHeat with Torch-DCT, which is implemented based on $torch.fft$. Our implemented vHeat-B (661 img/s) is much faster than Torch-DCT (367 img/s), validating that our implemented GPU-friendly matrix multiplication is significantly efficient.
@@ Appendix C Experimental Settings
Model configurations. The configurations of vHeat-T/S/B models are shown in Table  8. The FLOPs and training parameters are reported after reparameterization in HCOs.
@@
Layers
(2,2,6,2) (classification)
(2,2,5,2) (others)
(2,2,18,2) (classification)
(2,2,16,2) (others)
(2,2,18,2) (segmentation)
(4,4,20,4) (others)
Channels (96,192,384,768) (96,192,384,768)
(128,256,512,1024) (segmentation)
(96,192,384,768) (others)
Image Classification. Following the standard evaluation protocol used in [38], all vHeat series are trained from scratch for $300$ epochs and warmed up for the first $20$ epochs. We utilize the AdamW optimizer [40] during the training process with betas set to $(0.9,0.999)$, a momentum of $0.9$, a cosine decay learning rate scheduler, an initial learning rate of $2\times 10^{-3}$, a weight decay of $0.08$, and a batch size of $2048$. The drop path rates are set to $0.1/0.3/0.5$ for vHeat-T/S/B, respectively. Other techniques such as label smoothing ($0.1$) and exponential moving average (EMA) are also applied. No further training techniques are employed beyond these for a fair comparison. The training of vHeat-T/S/B takes 4.5/7/8.5 minutes per epoch on Tesla 16$\times$×V100 GPUs.
@@
Appendix D Additional Ablation Studies
@@
D.1 Interpolation of FVEs/$k$ for downstream tasks
We have tried several approaches to align the shape for ablation. (1) Directly interpolate FVEs to the target shape of the input image. (2) Add 0 to the lower right region of FVEs to align the target shape. (3) Add 0 to the lower right region of FVEs to $512\times 512$, and interpolate to the target shape. (4) Directly interpolate the predicted thermal diffusivity $k$ to the target shape. The results are summarized in Table  9. Through the comparison, we select adding 0, then interpolating FVEs to the target shape for all downstream tasks.
D.2 Plain vHeat model
We’ve tested the performance of plain vHeat-B on ImageNet-1K classification. Keeping the same as DeiT-B, plain vHeat-B has 12 HCO layers, 768 embedding channels and the patch size is set to 16. Results are shown in Table  10. The superiority of plain vHeat-B over DeiT-B also validates the effectiveness of vHeat model.
Table 10: Plain vHeat-B vs. DeiT-B on ImageNet-1K with 300 epochs supervised training.
@@ D.2 Plain vHeat model
D.3 Depth-wise convolution
We conduct experiments to validate the performance improvement from DWConv. We replace depth-wise convolution with layer normalization for vHeat-B. Results are summarized in Table  11, and vHeat-B achieves 83.8% Top-1 accuracy on ImageNet-1K classification, 0.2% lower than with DWConv, which validates the main gains come from the proposed HCO. Besides, when $k$ is fixed as a large value, e.g. $k=10.0$, replacing DWConv with layer normalization causes a significant performance drop (-0.7% top-1 accuracy). The comparison validates predicting $k$ by FVEs can effectively improve the robustness of vHeat.
@@ D.3 Depth-wise convolution
vHeat-B 84.0
vHeat-B 83.8 (-0.2)
vHeat-B (fix $k$=10.0) 83.6
vHeat-B (fix $k$=10.0) 82.9 (-0.7)
Appendix E Receptive Field Visualization
The Effective Receptive Field (ERF) [42] of an output unit denotes the region of input that contains elements with a non-negligible influence on that unit. In Fig. 7, ResNet, ConNeXT, and Swin have local ERF. DeiT [61] and vHeat exhibit global ERFs. The difference lies in that DeiT has a $\mathcal{O}(N^{2})$ complexity while vHeat enjoys $\mathcal{O}(N^{1.5})$ complexity.
Figure 7: Visualization of the effective receptive fields (ERF) [42]. The visualization of baseline models are provided from VMamba [36]. Pixels of higher intensity indicate larger responses with the central pixel.
@@ Appendix F Heat Conduction Visualization
Appendix G Analysis of $k$ in each layer
We calculate average values of $k$ in each layer of ImageNet-1K classification pre-trained vHeat-Tiny, Fig.  9. In stage 2 and stage 3, average values of $k$ corresponding to deeper layers are larger, indicating that the visual heat conduction effect of deeper layers is stronger, leading to faster and farther overall content propagation.
Figure 9: Average values of $k$ in each layer.
Appendix H Feature Map Visualization
We visualize the feature before/after HCO in a random layer in stage 2 with randomly selected images as input, Fig.  10. Before HCO, only a few regions of the foreground object are activated. After HCO, almost the entire foreground object is activated intensively.
@@ Appendix H Feature Map Visualization
Appendix I Ablation of the Linear+SiLU branch
The Linear + SiLU branch is a gated MLP unit, which is inspired by the design used in LLaMA [62]. We conducted ablation experiments, and the results, as shown in Table 12, demonstrate that the primary performance improvement of vHeat comes from the HCO, rather than the gated MLP.
Model #Param. FLOPs Top-1 acc. (%)
vHeat-B 68M 11.2G 84.0
Linear + ReLU 68M 11.2G 83.9
w/o (Linear + SiLU) 62M 10.3G 83.6
w/o HCO 49M 8.0G 76.7
Table 12: Ablation study of the Linear + SiLU branch.
Appendix J Comparison with SOTAs
We compare vHeat-B with other base-level SOTA visual representation models (MetaFormer-v2-M48, CAFormer-B36 [75], iFormer-B [51], and BiFormer-B [81]) on Top-1 Accuracy on ImageNet-1K and test throughput (an A100 GPU with 128 batch size), Figure 11. Our proposed vHeat demonstrates comparable performance with substantially improved test throughput.
Figure 11: Comparison of vHeat-B and other base-level SOTA vision models.